# An automated occultation network for gravitational mapping of the trans-neptunian solar system

### v1

@@ Abstract
We explore the potential of an array of $\mathcal{O}(100)$ small fixed telescopes, aligned along a meridian and automated to measure millions of occultations of Gaia stars by minor planets, to constrain gravitational signatures from a “Planet X” mass in the outer solar system. The accuracy of center-of-mass tracking for the occulters is limited by photon noise, uncertainties in asteroid shapes, and Gaia’s astrometry of the occulted stars. Using both parametric calculations and survey simulations, we assess the total information obtainable from occultation measurements of main-belt asteroids (MBAs), Jovian Trojans and trans-Neptunian objects (TNOs). We find that MBAs are the optimal target population due to their higher occultation rates and abundance of objects above LSST detection thresholds. A 10-year survey of occultations by MBAs and Trojans using an array of 200 40 cm telescopes at 5 km separation would achieve $5\sigma$ sensitivity to the gravitational tidal field of a $5M_{\oplus}$ Planet X at 800 AU for $>90\%$ of potential sky locations. This configuration corresponds to an initial cost of \$$\approx\$15$ million. While the occultation survey’s sensitivity to tidal forces improves rapidly with increasing number of telescopes, sensitivity to a Planet X becomes limited by degeneracy with the uncertain masses of large moonless TNOs. The 200-telescope survey would additionally detect $\approx 1800$ TNO occultations, providing detailed shape, size, and albedo information. The survey would also measure the Yarkovski effect on many individual MBAs, measure the masses of many asteroids involved in mutual gravitational deflections, and enable better searches for primordial black holes and departures from General Relativity.
@@ 1 Introduction
Main-belt asteroids (MBAs) and Jovian Trojans are more accessible targets, since they are closer to us and have larger apparent paths in the sky. These populations are also of interest in outer solar system studies, particularly due to their potential use as test particles in the search for distant gravitational signals. Rice & Laughlin (2019) suggested a network of $\sim$ 2000 telescopes spread over continental United States measuring occultations of Jovian Trojans, with the goal of detecting perturbations by a putative Planet X. Inspired by this idea, in a previous study (Gomes et al., 2023), we explored the ability of current astrometric data to constrain the presence of a distant tidal field, finding that LSST astrometry of Jovian Trojans does not significantly improve Planet X mass uncertainties. The possible contribution of an occultation array, however, remained unexplored.
@@ 2 Hardware characteristics and costs
Other important characteristics of the telescopes include the overall quantum efficiency $\eta$ of the atmosphere/telescope/aperture/detector combination, for which we assume 50%. The latitude, maximum observable airmass $X_{\rm max},$ and duty cycle $f_{\rm duty}$ are important to determining the yield of the occultations, where the latter is the fraction of occultations that pass over (or under!) the array that can actually be observed. We will default these to ±$\pm 30\arcdeg,$ 2.3, and 0.21, respectively. For an observatory at latitude ±$\pm 30\arcdeg,$ $f_{\rm duty}$ is limited to $<0.32$ by the fraction of near-ecliptic sky that is visible at airmass $<X_{\rm max}=2.3$ We reduce this by a factor $0.65$ to allow for losses due to clouds, down time, and bright twilight or moonlight that might be present—our simulations described in Section 4 will explicitly check each potential event against solar, lunar, and weather conditions rather than assuming an $f_{\rm duty}$ value.
@@ 2.1 Cost equation
  • Detectors: A CMOS or CCD “finder” with $\approx 10\arcmin$ FOV would not require particularly large format or high performance, so COTS hardware would suffice. Commercial single-channel photon-counting devices should work well too; we want to push for high QE, but the timing requirements (millisecond levels) and count rates (100–1000 Hz) that we require are more relaxed than typical specifications. The two detectors might cost \$$\approx\$15$k per station total.
@@
  • Least predictable would be the costs of negotiating and acquiring the leases for the sites. We ignore this important element for now.
In sum, a crude guess is hence $C_{s}=\$40$k and $C_{t}=\$60$K, leading to a \$10M cost for a 100-station array. We have neglected design costs (including software), which probably are independent of the array specifications, and maintenance/personnel costs, which might be expected to scale with the capital costs.
@@ 3.1 Test particle populations
We examine as target populations each of the three most populous reservoirs in the Solar System: the main-belt asteroids (MBAs), the Jovian Trojan asteroids, and the trans-Neptunian objects (TNOs). In this section we consider members of each population to share a common semi-major axis $a_{T}=2.6,5.2,$ and $42$ AU for the MBAs, Trojans, and TNOs, respectively. Aside from $a_{T},$ the other critical characteristic of the target population is $N_{T}(>d_{T}),$ the number of members larger than diameter $d_{T}.$ Figure 1 plots the values we have assumed for each population. Simple power laws are good fits to the regime where the MBAs and Trojans cataloged by the MPC [^2] https://www.minorplanetcenter.net/data are largely complete. For fainter MBAs, we adopt the formula used in the initial LSST predictions (LSST Science Collaboration et al., 2009), and for Trojans we extrapolate the power law fit. For TNOs, we adopt the rolling power law for $dN/dH$ found by Bernardinelli et al. (in prep.) to fit the $8.26<H_{r}<$ 8.2 range. We forecast that the known populations in future years will be dominated by objects detected by LSST. The expected $5\sigma$ detection levels for single-epoch LSST exposures in nominal conditions (dark, zenith, median seeing) [^3] https://pstn-054.lsst.io/ are equivalent to $m_{V}=24.4$ mag, which we hence consider a bound on the usable occulter population. There are $>10$ times more usable MBAs than Trojans, and potentially $10^{2}$ more MBAs than usable TNOs—but the physical displacements induced on a target by a tidal force scale as $a_{T}^{4}$ (for fixed number of orbits), so it is not a priori obvious which population will be most informative.
@@ 3.2 Measurement errors
The ultimate limit to the determination of orbits through occultations is the shape noise, our ignorance of the target’s projected center of mass location relative to the measured chords. Even perfect knowledge of the silhouette of the target would yield an uncertain center-of-mass (CM) estimate because of ignorance of the depth dimension, and variations in internal density cause additional CM shifts relative to the geometric centroid. In Section 4.2.1 we justify an estimate that $\sigma_{\rm SN}S=\alpha_{S}d_{T},$ with a nominal $\alpha_{S}=0.05$ giving the RMS ratio of the along-track CM distance from the center of a single chord to the mean radius of the target. The cross-track uncertainty from a single chord CM is larger, with $\sigma\approx{\rm min}(d_{T},D)/\sqrt{12},$ since a simple estimator is that the CM has a uniform probability of being somewhere in a region of width ${\rm min}(d_{T},D)$ centered on the observed chord. In this Section we will examine only the along-track information; the full simulation will use cross-track and multi-chord information as well.
@@
Appendix B derives the uncertainty $\sigma_{\gamma}$ in the center of a chord due to Poisson noise in the arrival of photons from both the occulted star (having unocculted photodetection rate $n_{\star}$) and some foreground/background count rate $n_{b}=Bn_{\star}$ due to atmospheric and zodiacal foreground signals, as well as sunlight light reflected from the occulting asteroid. For $B\rightarrow 0,$ bright stars, and large occulters, this uncertainty becomes $\sigma_{\gamma}=F/\sqrt{2n_{\star}t_{F}},$ where $F=\sqrt{a_{T}\lambda_{0}/2}$ is the Fresnel scale which characterizes the distance over which the occulter’s shadow transitions from fully bright to fully dark. This transition region takes a time $t_{F}=F/v_{\perp}$ to sweep over the telescope, where $v_{\perp}$ is the relative speed of the occulter and the observer transverse to the line of sight. The Fresnel scales $F$ and $t_{F}$ are $\approx 340$ m and 11 ms, respectively, for typical MBAs, growing as $\sqrt{a_{T}}$ for more distant occulter samples.
@@
Figure 2: Sources of uncertainty in the along-track asteroid position from a single occultation chord are plotted vs magnitude of the occulted star. At left is for an occulter diameter $d_{T}=3$ km, at right $d_{T}=0.5$ km, both sources at distance 2.6 AU. Each set of lines shows values of the photon noise $\sigma_{\gamma}$ for new, quarter, and full moons, and three different telescope diameters $d$ are plotted in each case. The magenta curve plots the expected final Gaia positional uncertainties $\sigma_{G}$, and the horizontal line marked “shape noise” marks the floor $\sigma_{\rm SN}$ set by the unknown shape and density distribution of the occulter. The “detection threshold” marks the uncertainty $\sigma_{\gamma}=d_{T}/4$ that we consider an upper limit for detectable occultations, and the $\sigma_{\gamma}$ curves terminate at this level. The Fresnel scale is also plotted, and the small triangles mark the $G$ magnitudes at which background counts equal star counts for different moon phases. For the 0.5 m occulter, near the diffraction scale, almost all occultations are limited by photon noise, and the sky brightness is unimportant. For the larger occulter, fainter Gaia stars yield detectable occultations, and the shape noise limits the measurement accuracy for much of the detectable regime.
@@ 3.3 Event rates
The second line makes use of the following estimates: first, we estimate the mean sky density of useful Gaia stars within ±$\pm 10\arcdeg$ of the ecliptic as $n_{G}(<m_{G})\approx 3500\times 10^{0.3(m_{G}-18)}\,\text{deg}^{-2}.$ This estimate does not count as “useful” those Gaia stars in crowded low-Galactic-latitude regions.
@@
Not all positional information on a target can be used to constrain unexpected gravitational accelerations. There are six degrees of freedom in each target’s initial state vector, which must also be solved using the occultation data. Previous data (nominally from LSST) would be available too, but will probably be 10$\gtrsim$ 10$\times$ 10 × less accurate than the occultation data and hence not useful at occulation-survey accuracy—otherwise we would not be doing an occultation survey! We therefore assume that, for a given asteroid, the three most precise occultation measurements will be used to constrain the initial state, and information on perturbations will begin to accrue only from the 4th most accurate occultation observation. Thus a target which does not produce $\geq 4$ measurable occultations does not contribute at all to the $\sigma_{\rm tot}$ figure of merit. When forecasting the actual number of occultations, we assume each target draws from a Poisson distribution with mean given by Equation (4), rather than assigning every asteroid the mean number of observed occultations.
@@ 3.4 Results: MBAs
Figure 3: These figures illustrate the performance predicted by the parametric model for an occultation array targeting MBAs using a configuration of $N=200$ telescopes of diameter $d=0.4$ m at $D=2$ km spacing, near the optimum for a \$15M nominal capital cost. The uncertainty on the RMS size of an anomalous gravitational displacement after a 10-year survey is just $\sigma_{\rm tot}=4.4$ cm. At left is the total measurement uncertainty in the occulter position for single-chord observations, vs. the occulter diameter and magnitude of the occulted star. Overlaid are several important dividing lines: asteroids below the $\rho=1$ line have radius below the Fresnel scale, only partially blocking stellar light. Those below the “Unknown” line are too faint to be found in the nominal LSST single-epoch data. Those to the right of the $B=0.5$ contour are substantially degraded by “background” light from the night sky and the asteroid’s reflected light. Those to the left of the $\sigma_{\gamma}=\sigma_{\rm SN}$ contour are limited by shape noise in determining the center of mass of the asteroid. Those to the right of the $\sigma_{\gamma}=d_{T}/4$ curve have too much photon noise for the occultation to be securely detected. The right-hand plot shows the density of information on gravitational perturbations in the $d_{T}-m_{G}$ plane, which is heavily concentrated in the region with the most numerous occultations—the asteroids near the Fresnel scale, with the faintest stars that produce detectable occultations. The $\langle N_{e}\rangle=4$ contour marks the magnitude of the star that will on average be the 4th-brightest occulted by an asteroid of a given size. Occulters with $N_{e}<4$ produce negligible information on gravitational perturbations.
@@
  • The tradeoff between telescope diameter $d_{T}$ and number of telescopes $N$ at fixed cost minimizes $\sigma_{\rm tot}$ when the cost of the telescope $C_{t}(d_{T}/0.5\,{\rm m})^{2}$ is 10–20% higher than the fixed station costs $C_{s}$. In other words: spread the funds over as many (smaller) telescopes as possible until the per-station costs are about to dominate. When $C_{s}=\$40{\rm k}$, this leads to optimal telescope size $d_{T}\approx 0.45$ m.

### v2

@@ Abstract
We explore the potential of an array of $\mathcal{O}(100)$ small fixed telescopes, aligned along a meridian and automated to measure millions of occultations of Gaia stars by minor planets, to constrain gravitational signatures from a “Planet X” mass in the outer solar system. The accuracy of center-of-mass tracking of occulters is limited by photon noise, uncertainties in asteroid shapes, and Gaia’s astrometry of the occulted stars. Using both parametric calculations and survey simulations, we assess the total information obtainable from occultation measurements of main-belt asteroids (MBAs), Jovian Trojans and trans-Neptunian objects (TNOs). We find that MBAs are the optimal target population due to their higher occultation rates and abundance of objects above LSST detection thresholds. A 10-year survey of occultations by MBAs and Trojans using an array of 200 40 cm telescopes at 5 km separation would achieve $5\sigma$ sensitivity to the gravitational tidal field of a $5M_{\oplus}$ Planet X at 800 AU for $>90\%$ of potential sky locations. This configuration corresponds to an initial cost of \$$\approx\$15$ million. While the survey’s sensitivity to tidal forces improves rapidly with increasing number of telescopes, sensitivity to a Planet X becomes limited by degeneracy with the uncertain masses of large moonless TNOs. The 200-telescope survey would additionally detect $\approx 1800$ TNO occultations, providing detailed shape, size, and albedo information. It would also measure the Yarkovsky effect on many individual MBAs, measure masses of many asteroids involved in mutual gravitational deflections, and enable better searches for primordial black holes and departures from General Relativity.
@@ 1 Introduction
Main-belt asteroids (MBAs) and Jovian Trojans are more accessible targets, since they are closer to us and have larger apparent paths in the sky. These populations are also of interest in outer solar system studies, particularly due to their potential use as test particles in the search for distant gravitational signals, which could provide constraints on different “unseen planet” hypotheses (generally referred to as Planet X), such as the Planet Nine hypothesis evoked to explain a perceived orbital clustering of TNOs  (Trujillo & Sheppard, 2014; Sheppard & Trujillo, 2016; Batygin & Brown, 2016; Brown & Batygin, 2016). Rice & Laughlin (2019) suggested a network of $\sim$ 2000 telescopes spread over continental United States measuring occultations of Jovian Trojans, with the goal of detecting perturbations by a putative Planet X. Inspired by this idea, in a previous study (Gomes et al., 2023), we explored the ability of current and planned LSST astrometric data to constrain the presence of a distant tidal field, finding that LSST astrometry of Jovian Trojans does not significantly improve Planet X mass uncertainties. The possible contribution of an occultation array, however, remained unexplored beyond Rice and Laughlin’s initial foray.
@@
This paper will focus on optimizing a telescope array for minor-body astrometry and tidal-field detection, but an array of $\gtrsim 100$ agile telescopes could also advance many other science goals. The astrometric data would also improve the solar system ephemeris generally, constrain the flux of asteroid-mass primordial black holes (Tran et al., 2024), and vastly improve knowledge of the sizes of asteroids and their non-gravitational forces that drive much migration. A review of knowledge gained on TNOs through occultations is given by Sicardy et al. (2024), and in Section 5 we show that such an array would measure the shapes and albedos of 100$\approx$ 100$\times$ 100 × more TNOs than have been characterized to date. Individual stations would be available for other observations when they are not active for an occultation, enabling high-duty-cycle monitoring for asteroid impacts, supernova shock breakouts, and many other transient or variable phenomena occurring on time scales or cadences inaccessible to LSST.
@@ 2 Hardware characteristics and costs
Other important characteristics of the telescopes include the overall quantum efficiency $\eta$ of the atmosphere/telescope/aperture/detector combination, for which we assume 50%. The latitude, maximum observable airmass $X_{\rm max},$ and duty cycle $f_{\rm duty}$ are important to determining the yield of the occultations, where the latter is the fraction of occultations that pass over (or under) the array that can actually be observed. We will default these to ±$\pm 30\arcdeg,$ 2.3, and 0.21, respectively. For an observatory at latitude ±$\pm 30\arcdeg,$ $f_{\rm duty}$ is limited to $<0.32$ by the fraction of near-ecliptic sky that is visible at airmass $<X_{\rm max}=2.3$ We reduce this by a factor $0.65$ to allow for losses due to clouds, down time, and bright twilight or moonlight that might be present—our simulations described in Section 4 will explicitly check each potential event against solar, lunar, and weather conditions rather than assuming an $f_{\rm duty}$ value.
@@ 2.1 Cost equation
  • Detectors: A CMOS or CCD “finder” with $\approx 10\arcmin$ FOV would not require particularly large format or high performance, so commercial off-the-shelf hardware would suffice. Commercial single-channel photon-counting devices should work well too; we want to push for high QE, but the timing requirements (millisecond levels) and count rates (100–1000 Hz) that we require are more relaxed than typical specifications. The two detectors might cost \$$\approx\$15$k per station total.
@@
  • Least predictable would be the costs of negotiating and acquiring the leases for the sites. We ignore this important element for now, because any believable estimate would depend heavily on choice of site and require research well outside of our astronomical expertise.
In sum, a crude guess is hence $C_{s}=\$40$k and $C_{t}=\$60$K, leading to a \$10M cost for a 100-station array. We have neglected design costs (including software), which probably are independent of the array specifications. The above costs are all for capital outlays. There would of course be ongoing operating costs for maintenance/personnel. At most observatories these are 5–10% of capital cost annually, accumulating to be similar to capital costs over the project lifetime. However, this unusual observatory might be more expensive because of its geographically distributed nature, or less expensive because its replication of components makes maintainance more efficient. An engineering and cost study is needed.
@@ 3.1 Test particle populations
We examine as target populations each of the three most populous reservoirs in the Solar System: the main-belt asteroids (MBAs), the Jovian Trojan asteroids, and the trans-Neptunian objects (TNOs). In this section we consider members of each population to share a common semi-major axis $a_{T}=2.6,5.2,$ and $42$ AU for the MBAs, Trojans, and TNOs, respectively. Aside from $a_{T},$ the other critical characteristic of the target population is $N_{T}(>d_{T}),$ the number of members larger than diameter $d_{T}.$ Figure 1 plots the values we have assumed for each population. Simple power laws are good fits to the regime where the MBAs and Trojans cataloged by the MPC [^2] https://www.minorplanetcenter.net/data are largely complete. For fainter MBAs, we adopt the formula used in the initial LSST predictions (LSST Science Collaboration et al., 2009), and for Trojans we extrapolate the power law fit. For TNOs, we adopt the rolling power law for $dN/dH$ found by Bernardinelli et al. (2025) to fit the $8.26<H_{r}<$ 8.2 range. We forecast that the known populations in future years will be dominated by objects detected by LSST. The expected $5\sigma$ detection levels for single-epoch LSST exposures in nominal conditions (dark, zenith, median seeing) [^3] https://pstn-054.lsst.io/ are equivalent to $m_{V}=24.4$ mag, which we hence consider a bound on the usable occulter population. There are $>10$ times more usable MBAs than Trojans, and potentially $10^{2}$ more MBAs than usable TNOs—but the physical displacements induced on a target by a tidal force scale as $a_{T}^{4}$ (for fixed number of orbits), so it is not a priori obvious which population will be most informative.
@@ 3.2 Measurement errors
The ultimate limit to the determination of orbits through occultations is the shape noise, our ignorance of the target’s projected center of mass location relative to the measured chords. Even perfect knowledge of the silhouette of the target would yield an uncertain center-of-mass (CM) estimate because of ignorance of the depth dimension, and variations in internal density cause additional CM shifts relative to the geometric centroid. In Section 4.2.1 we justify an estimate that $\sigma_{\rm SN}=\alpha_{S}d_{T},$ with a nominal $\alpha_{S}=0.05$ giving the RMS ratio of the along-track CM distance from the center of a single chord to the mean radius of the target. The cross-track uncertainty from a single chord CM is larger, with $\sigma\approx{\rm min}(d_{T},D)/\sqrt{12},$ since a simple estimator is that the CM has a uniform probability of being somewhere in a region of width ${\rm min}(d_{T},D)$ centered on the observed chord. In this Section we will examine only the along-track information; the full simulation will use cross-track and multi-chord information as well.
@@
Appendix B derives the uncertainty $\sigma_{\gamma}$ in the center of a chord due to Poisson noise in the arrival of photons from both the occulted star (having unocculted photodetection rate $n_{\star}$) and some foreground/background count rate $n_{b}$ due to atmospheric and zodiacal foreground signals, as well as sunlight reflected from the occulting asteroid. The background ratio is defined as $B=n_{b}/n_{\star}$. For $B\rightarrow 0,$ bright stars, and large occulters, this uncertainty becomes $\sigma_{\gamma}=F/\sqrt{2n_{\star}t_{F}},$ where $F=\sqrt{a_{T}\lambda_{0}/2}$ is the Fresnel scale which characterizes the distance over which the occulter’s shadow transitions from fully bright to fully dark. This transition region takes a time $t_{F}=F/v_{\perp}$ to sweep over the telescope, where $v_{\perp}$ is the relative speed of the occulter and the observer transverse to the line of sight. The Fresnel scales $F$ and $t_{F}$ are $\approx 340$ m and 11 ms, respectively, for typical MBAs, growing as $\sqrt{a_{T}}$ for more distant occulter samples.
@@
Figure 2: Sources of uncertainty in the along-track asteroid position from a single occultation chord are plotted vs magnitude of the occulted star. At left is for an occulter diameter $d_{T}=3$ km, at right $d_{T}=0.5$ km, both sources at distance 2.6 AU. Each set of lines shows values of the photon noise $\sigma_{\gamma}$ for new, quarter, and full Moons, and three different telescope diameters $d$ are plotted in each case. The magenta curve plots the expected final Gaia positional uncertainties $\sigma_{G}$, and the horizontal line marked “shape noise” marks the floor $\sigma_{\rm SN}$ set by the unknown shape and density distribution of the occulter. The “detection threshold” marks the uncertainty $\sigma_{\gamma}=d_{T}/4$ that we consider an upper limit for detectable occultations, and the $\sigma_{\gamma}$ curves terminate at this level. The Fresnel scale is also plotted, and the small triangles mark the $G$ magnitudes at which background counts equal star counts for different moon phases. For the 0.5 km occulter, near the diffraction scale, almost all occultations are limited by photon noise, and the sky brightness is unimportant. For the larger occulter, fainter Gaia stars yield detectable occultations, and the shape noise limits the measurement accuracy for much of the detectable regime.
@@ 3.3 Event rates
The final expression makes use of the following estimates: first, we estimate the mean sky density of useful Gaia stars within ±$\pm 10\arcdeg$ of the ecliptic as $n_{G}(<m_{G})\approx 3500\times 10^{0.3(m_{G}-18)}\,\text{deg}^{-2}.$ This estimate does not count as “useful” those Gaia stars in crowded low-Galactic-latitude regions.
@@
Not all positional information on a target can be used to constrain unexpected gravitational accelerations. There are six degrees of freedom in each target’s initial state vector, which must also be solved using the occultation data. Previous data (nominally from LSST) would be available too, but will probably be 10$\gtrsim$ 10$\times$ 10 × less accurate than the occultation data and hence not useful at occultation-survey accuracy—otherwise we would not be doing an occultation survey! We therefore assume that, for a given asteroid, the three most precise occultation measurements will be used to constrain the initial state, and information on perturbations will begin to accrue only from the 4th most accurate occultation observation. Thus a target which does not produce $\geq 4$ measurable occultations does not contribute at all to the $\sigma_{\rm tot}$ figure of merit. When forecasting the actual number of occultations, we assume each target draws from a Poisson distribution with mean given by Equation (4), rather than assigning every asteroid the mean number of observed occultations.
@@ 3.4 Results: MBAs
Figure 3: These figures illustrate the performance predicted by the parametric model for an occultation array targeting MBAs using a configuration of $N=200$ telescopes of diameter $d=0.4$ m at $D=2$ km spacing, near the optimum for a \$15M nominal capital cost. The uncertainty on the RMS size of an anomalous gravitational displacement after a 10-year survey is just $\sigma_{\rm tot}=4.4$ cm. At left is the total measurement uncertainty in the occulter position for single-chord observations, vs. the occulter diameter and magnitude of the occulted star. Overlaid are several important dividing lines: asteroids below the $d_{t}=2F$ line have radius below the Fresnel scale, only partially blocking stellar light. Those below the “Unknown” line are too faint to be found in the nominal LSST single-epoch data. Those to the right of the $B=0.5$ contour are substantially degraded by “background” light from the night sky and the asteroid’s reflected light. Those to the left of the $\sigma_{\gamma}=\sigma_{\rm SN}$ contour are limited by shape noise in determining the center of mass of the asteroid. Those to the right of the $\sigma_{\gamma}=d_{T}/4$ curve have too much photon noise for the occultation to be securely detected. The right-hand plot shows the density of information on gravitational perturbations in the $d_{T}-m_{G}$ plane, which is heavily concentrated in the region with the most numerous occultations—the asteroids near the Fresnel scale, with the faintest stars that produce detectable occultations. The $\langle N_{e}\rangle=4$ contour marks the magnitude of the star that will on average be the 4th-brightest occulted by an asteroid of a given size. Occulters with $N_{e}<4$ produce negligible information on gravitational perturbations.
@@
  • The tradeoff between telescope diameter $d_{T}$ and number of telescopes $N$ at fixed cost minimizes $\sigma_{\rm tot}$ when the cost of the telescope $C_{t}(d_{T}/0.5\,{\rm m})^{2}$ is 10–20% higher than the fixed station costs $C_{s}$. In other words, onde should spread the funds over as many (smaller) telescopes as possible until the per-station costs are about to dominate. When $C_{s}=\$40{\rm k}$, this leads to optimal telescope size $d_{T}\approx 0.45$ m.